# Can Graph Neural Networks Learn Language with Extremely Weak Text Supervision?

### v2

@@ 5.5 Efficiency and Embedding Analysis
Without fine-tuning the GNN or LLM, the prompt-based methods have better parameter efficiency. As shown in Figure 5 (left), our ImprovedAIO and Morpher require similar numbers of parameters with AIO [51], which is $0.032\%$ to $0.46\%$ compared to either tune the LLM (RoBERTa) or GNN (GCN). Due to such parameter efficiency, our methods learn better graph representations given few-shot data. We visualize the graph embeddings of CiteSeer and MSRC_21C in Figure 4 and calculate the silhouette score, a metric for cluster quality ($\uparrow$) ranged in $[-1,1]$. It turns out that our multimodal prompting leads to better adaptation.
@@ 5.6 Hyperparameter and Ablation Study
We conduct the hyperparameter study by choosing and testing various numbers of graph prompt tokens for both ImprovedAIO and Morpher. The results are shown in Figure 5 (middle), from which we can observe that both methods are generally stable, and Morpher constantly outperforms ImprovedAIO under different choices. To verify the necessity of each component in our design, we compare Morpher and ImprovedAIO with multiple variants, respectively, and report the result in Figure 5 (right). We observe that removing any component would result in a performance drop. Additionally, our comparison of Morpher with ImprovedAIO throughout the experiments demonstrates that our multimodal design would lead to improvement over the uni-modal prompting of GNNs.
Table 4: Effectiveness (F1 score) of Morpher with ELECTRA [62] as the text encoder.
GNN pretraining MUTAG ENZYMES PROTEINS MSRC_21C
GraphCL + GCN 78.17 15.79 65.66 47.19
GraphCL + GAT 75.75 11.37 65.66 49.01
GraphCL + GT 77.04 14.68 72.70 44.09
SimGRACE + GCN 70.99 12.41 67.77 48.44
SimGRACE + GAT 77.51 13.31 67.78 49.43
SimGRACE + GT 73.55 15.76 70.28 44.50
In the main experiments, we use RoBERTa as Morpher’s text encoder. We also conduct experiments to verify the effectiveness of our proposed Morpher with ELECTRA [62] and DistilBERT [63] as the text encoder. Due to space limitation, we only show the F1 score of using ELECTRA in Table 4, and more detailed experiment data can be found in Appendix C.3. In general, using ELECTRA and DistilBERT results in similar performance compared to using RoBERTa, showing the robustness of Morpher with respect to the language encoder.
As for the robustness with respect to the pre-trained GNNs, in the main experiments, we adopt two pre-train methods, GraphCL and SimGRACE to pre-train three different GNN architectures: GCN, GAT and GT. We further conduct experiments using GNNs pre-trained from GraphMAE [71] and MVGRL [72]. Due to the space limitation, we report the results and discuss in Appendix C.4.
5.7 Morpher on MolecureNet with More Text Supervision
We demonstrate that, though not specifically designed for any downstream applications, the Morpher framework has the potential to be used in various tasks where there is more text supervision compared to previous experiments. As for a case study, We use bace (inhibitors of human beta-secretase), tox21 (toxicology in the 21st century) and hiv (inhibit HIV replication) from MolecureNet [74]. These three datasets have 1513, 7831, and 41127 graphs to classify, respectively. In these datasets, each graph label is associated with a text description. The tasks on bace and hiv are bio-activity prediction and the task on tox21 is toxicity prediction. To adopt Morpher, we use GraphCL to pre-train the GAT model and initialize the text prompts and text labels using those from GIMLET [75].
bace 0.5126 0.6656 0.5648 0.729 0.697 0.6231 0.6513 0.6858
tox21 0.4917 0.5757 0.4946 0.652 0.754 0.6769 0.7275 0.7459
hiv 0.6120 0.5026 0.3385 0.721 0.729 0.5742 0.7034 0.7283
KVPLM [76], MoMu [16], Galactica-1.3B [77] are zero-shot predictors for the three tasks; GIMLET-64M-50-shots is the GIMLET [75] model fine-tuned on 50 additional training samples333the performance of GIMLET and other baselines are directly from the GIMLET paper [75].; GAT-1M-fully-supervised uses all the training data to train a GAT. Our Morpher-k-shots uses only k training samples. From the results, first, using only 10 training samples, Morpher can outperform the zero-shot baselines KVPLM, MoMu, and Galactica-1.3B. Second, using only 50 shots, Morpher can achieve similar performance with the fully supervised GAT. Third, using the same amount of few-shot data (50 shots), Morpher-50 outperforms GIMLET-64M-50-shots on tox21 and hiv, the two largest datasets among the three. This means our graph-text multi-modal prompt learning, with much fewer learnable parameters ($\sim 50K$), is more sample-efficient than fine-tuning language model encoder.
@@ 6 Related Work
Recently, a surge of graph pre-training strategies have emerged to address the issue of label scarcity in graph representation learning [41, 78, 73, 79, 80, 81, 82]. The main idea of pre-trained graph models is to capture general graph information across different tasks and transfer this knowledge to the target task using techniques such as contrastive predictive coding [70, 83, 84, 85, 67], context prediction [86, 87], prompt tuning [73, 59], and mutual information maximization [86, 88, 89, 90, 91]. For instance, [41] proposes to learn transferable structural information from three levels of graph topology, including node-level, subgraph-level, and graph-level. Different from these approaches, this paper aims to build up foundational GNNs by leveraging multi-modal prompt learning techniques.
@@
Prompting is now mainstream for adapting NLP tasks, and recent studies exploring prompt learning for GNNs mark a thriving research area [92, 93]. It is a promising way to adapt GNNs to downstream tasks through token-level [59, 94, 95, 73, 96] or graph-level [51, 97, 98] prompting. Among all the existing methods, All-in-one (AIO) [51] is the only algorithm to learn tunable graph prompts for node-level, edge-level or graph-level downstream tasks given few-shot labeled data (Table 8). Based on our improved AIO, we present a pioneer study to explore learning prompts in multiple modalities simultaneously while keeping the pre-trained models frozen.
@@
Inspired by the advances of large language models in NLP [99], researchers have begun to explore their potential for graph-related tasks [100]. Current approaches can be divided into two main categories. The first category employs LLMs as pre-trained feature extractors to enhance GNNs [101, 102, 103]. For example, GLEM [104] proposes to input the language representation as initial features for the GNN and train them iteratively. The second category focuses on integrating graph structures directly into LLM architectures [105, 106, 107]. A notable example is Patton [108], which pre-trains a joint architecture on text-attributed graphs. Despite these advancements, none of them have explored the collaboration between LLMs and GNNs under graph prompt learning.
@@ 7 Conclusion
In this work, we introduce Morpher, the first multimodal prompt learning paradigm that can semantically adapt pre-trained GNNs to downstream tasks with the help of LLM, while keeping both the pre-trained models frozen. To build Morpher, we first analyze the limitations of the state-of-the-art graph prompting technique and propose an improved version. Through extensive experiments, we demonstrate that our improved AIO can achieve outperformance, and our Morpher has further improvements in few-shot, multi-level task, or domain transfer settings. Additionally, using Morpher, we build the first GNN zero-shot classifier prototype that can be generalized to novel testing classes.
Appendix A Proof of Lemma 3.1
Proof.
The cross-entropy loss between the true distribution $p(\cdot)$ and the predicted distribution $q(\cdot)$ is given by:
$\text{CE}(p,q)$ $=-\sum_{y}p(y)\log q(y)$
where $q(y)=Pr(c(\mathbf{x})=y)$.
To find the optimal classification, we minimize the cross-entropy loss subject to the constraint $\sum_{y}q(y)=1$. We define the Lagrangian as:
$\mathcal{L}(q,\lambda)$ $=-\sum_{y}p(y)\log q(y)+\lambda\left(\sum_{y}q(y)-1\right)$
For any $y\in\mathcal{Y}$, take the derivative of $\mathcal{L}$ with respect to $q(y)$ and $\lambda$ and set them to zero, we get:
$\frac{\partial\mathcal{L}}{\partial q(y)}$ $=-\frac{p(y)}{q(y)}+\lambda=0$
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=\sum_{y}q(y)-1=0$
Solving these equations, we find:
$q(y)$ $=\frac{p(y)}{\lambda}$
$\sum_{y}q(y)$ $=\sum_{y}\frac{p(y)}{\lambda}=\frac{1}{\lambda}\sum_{y}p(y)=1$
Therefore, $\lambda=1$ and $q(y)=p(y)$.
Thus, the optimal classification is $Pr(c(\mathbf{x})=y)=p(y)$.
Appendix B Dataset Details
B.1 Dataset Statistics
@@
B.2 Text Labels
@@
GPF-Plus [59] token-level $\surd$ $\times$ $\times$ $\surd$ $\times$
Gprompt [60] token-level $\surd$ $\times$ $\surd$ $\surd$ $\times$
VNT [94] token-level $\times$ $\times$ $\surd$ $\surd$ $\times$
ULTRA-DP [95] token-level $\surd$ $\times$ $\times$ $\surd$ $\times$
GPPT [73] token-level $\surd$ $\times$ $\times$ $\surd$ $\times$
SGL-PT [96] token-level $\surd$ $\times$ $\times$ $\surd$ $\times$
SAP [98] graph-level $\surd$ $\times$ $\surd$ $\surd$ $\times$
PRODIGY [97] graph-level $\surd$ $\surd$ $\surd$ $\times$ $\times$
All-in-one (AIO) [51] graph-level $\surd$ $\surd$ $\surd$ $\surd$ $\times$
@@
Appendix C Experiment Details
C.1 Reproducibility
@@ C.2 Implementation Details
C.3 Experiment with ELECTRA and DistilBERT
On the LLM pre-training side, RoBERTa is one of the most advanced encoder-only LLMs until now, and we have demonstrated the effectiveness with RoBERTa serving on the LLM side in the Morpher paradigm. Additionally, we conducted experiments with ELECTRA [62] and DistilBERT [63]. Using these two LLMs, Morpher can also achieve comparable performances to RoBERTa. The results are shown as follows.
Table 9: Few-shot graph classification performance (%) of Morpher with ELECTRA [62] as language encoder. Other experiment settings are identical to the main experiment.
@@ C.3 Experiment with ELECTRA and DistilBERT
Table 10: Few-shot graph classification performance (%) of Morpher with DistilBERT [63] as language encoder. Other experiment settings are identical to the main experiment.
@@
C.4 Experiment with GNNs trained using GraphMAE and MVGRL
In the main pages, we used GraphCL and SimGRACE to show that Morpher achieves better performance given a pre-trained GNN. Additionally, to further verify the robustness of Morpher over the pre-train method, we conducted experiments on the pre-trained GNNs using GraphMAE [71] and MVGRL [72]. We use GCN as the GNN backbone and RoBERTa as the LLM encoder, and the results are reported as follows.
Table 11: Few-shot graph classification performance (%) of Morpher with the GNN pre-trained by GraphMAE [71]. Other experiment settings are identical to the main experiment.
@@ C.4 Experiment with GNNs trained using GraphMAE and MVGRL
Table 12: Few-shot graph classification performance (%) of Morpher with the GNN pre-trained by MVGRL [72]. Other experiment settings are identical to the main experiment.

### v3

@@ 5.5 Efficiency and Embedding Analysis
Without fine-tuning the GNN or LLM, the prompt-based methods have better parameter efficiency. As shown in Figure 4 (left), our ImprovedAIO and Morpher require similar numbers of parameters with AIO (Sun et al., 2023a), which is $0.032\%$ to $0.46\%$ compared to either tune the LLM (RoBERTa) or GNN (GCN). Due to such parameter efficiency, our methods learn better graph representations given few-shot data. We visualize the graph embeddings of CiteSeer and MSRC_21C in Figure 3 and calculate the silhouette score, a metric for cluster quality ($\uparrow$) ranged in $[-1,1]$. It turns out that our Morpher leads to better adaptation.
@@ 5.6 Hyperparameter and Ablation Study
We conduct the hyperparameter study by choosing and testing various numbers of graph prompt tokens for both ImprovedAIO and Morpher. The results are shown in Figure 4 (middle), from which we can observe that both methods are generally stable, and Morpher constantly outperforms ImprovedAIO under different choices. To verify the necessity of each component in our design, we compare Morpher and ImprovedAIO with multiple variants, respectively, and report the result in Figure 4 (right). We observe that removing any component would result in a performance drop.
We also conduct experiments to verify the effectiveness of our proposed Morpher with ELECTRA (Clark et al., 2020) and DistilBERT (Sanh et al., 2019) as the text encoder in Appendix B.3 due to space limitation. In general, Morpher is robust with respect to the language encoder. As for the robustness with respect to the pre-trained GNNs, we further conduct experiments using GNNs pre-trained from GraphMAE (Hou et al., 2022) and MVGRL (Hassani and Ahmadi, 2020). Due to the space limitation, the results are in Appendix B.4.
@@ 6 Related Work
Recently, a surge of graph pre-training strategies have emerged (Hu et al., 2020a; Lu et al., 2021; Jing et al., 2021; Zhou et al., 2022a). The main idea of pre-trained graph models is to capture general graph information across different tasks and transfer this knowledge to the target task using techniques such as contrastive predictive coding (Khosla et al., 2020; Xia et al., 2022), context prediction (Hu et al., 2020b), and mutual information maximization (Sun et al., 2020). Different from these approaches, this paper aims to adapt pre-trained GNNs by leveraging multi-modal prompt learning techniques.
@@
Recent studies exploring prompt learning for GNNs mark a thriving research area (Sun et al., 2023b; Wu et al., 2023b). It is a promising way to adapt GNNs to downstream tasks through token-level (Fang et al., 2023; Tan et al., 2023; Chen et al., 2023a; Sun et al., 2022; Zhu et al., 2023) or graph-level (Sun et al., 2023a; Huang et al., 2023; Ge et al., 2023) prompting. Among all the existing methods, All-in-one (AIO) (Sun et al., 2023a) is the only algorithm to learn tunable graph prompts for multi-level downstream tasks (Table 6). Based on our improved AIO, we present a pioneer study to explore learning prompts in multiple modalities simultaneously while keeping the pre-trained models frozen.
@@
LLMs’ potential for graph-related tasks (Jin et al., 2023a) has been explored recently. The first category employs LLMs as pre-trained feature extractors to enhance GNNs (Duan et al., 2023; Chien et al., 2021; Zhu et al., 2021). The second category focuses on integrating graph structures directly into LLM architectures (Yang et al., 2021; Zhang et al., 2022b; Jin et al., 2023d, c). Despite these advancements, none of them have explored the collaboration between LLMs and GNNs with extremely weak text supervision and under graph prompt learning.
@@ 7 Conclusion
We present Morpher, the first multimodal prompt learning paradigm leveraging LLMs to semantically adapt pre-trained GNNs to downstream tasks with extremely weak text supervision. Addressing limitations of existing graph prompting techniques, we demonstrate through extensive experiments that Morpher excels in few-shot, multi-level tasks, and domain transfer. Notably, Morpher enables generalization to novel testing classes.
Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation Award No. IIS-19-56151 and Award No. IIS-2117902, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329, and IBM-Illinois Discovery Accelerator Institute - a new model of an academic-industry partnership designed to increase access to technology education and skill development to spur breakthroughs in emerging areas of technology. The views and conclusions are those of the authors and should not be interpreted as representing the official policies of the funding agencies or the government.
8 Limitations
Our Multimodal graph prompt learning paradigm assumes the “pre-train + prompt” framework to learn transferable graph representations, yet there could be other paths to achieve graph-related foundation models. Also, graph prompt learning only works on the graph neural network architecture, and might not work for other architectures that are proposed in the future. Another limitation of this work is the requirement of language encoder. While RoBERTa is one of the most advanced encoder-only language models and can be considered an LLM with over 0.1B parameters, more recent LLMs such as Llama or Mistral cannot be used in Morpher because they are decoder-only LLMs and do not explicitly have an encoder.
Appendix A Dataset Details
@@
A.1 Dataset Statistics
A.2 Text Labels
@@
GPF-Plus (Fang et al., 2023) token-level $\surd$ $\times$ $\times$ $\surd$ $\times$
Gprompt (Liu et al., 2023d) token-level $\surd$ $\times$ $\surd$ $\surd$ $\times$
VNT (Tan et al., 2023) token-level $\times$ $\times$ $\surd$ $\surd$ $\times$
ULTRA-DP (Chen et al., 2023a) token-level $\surd$ $\times$ $\times$ $\surd$ $\times$
GPPT (Sun et al., 2022) token-level $\surd$ $\times$ $\times$ $\surd$ $\times$
SGL-PT (Zhu et al., 2023) token-level $\surd$ $\times$ $\times$ $\surd$ $\times$
SAP (Ge et al., 2023) graph-level $\surd$ $\times$ $\surd$ $\surd$ $\times$
PRODIGY (Huang et al., 2023) graph-level $\surd$ $\surd$ $\surd$ $\times$ $\times$
All-in-one (AIO) (Sun et al., 2023a) graph-level $\surd$ $\surd$ $\surd$ $\surd$ $\times$
@@
Appendix B Experiment Details
B.1 Reproducibility
@@ B.2 Implementation Details
B.3 Experiment with ELECTRA and DistilBERT
On the LLM pre-training side, RoBERTa is one of the most advanced encoder-only LLMs until now, and we have demonstrated the effectiveness with RoBERTa serving on the LLM side in the Morpher paradigm. Additionally, we conducted experiments with ELECTRA (Clark et al., 2020) and DistilBERT (Sanh et al., 2019). Using these two LLMs, Morpher can also achieve comparable performances to RoBERTa. The results are shown in Table 8 and in Table 7.
Table 7: Few-shot graph classification performance (%) of Morpher with ELECTRA (Clark et al., 2020) as language encoder. Other experiment settings are identical to the main experiment.
@@ B.3 Experiment with ELECTRA and DistilBERT
Table 8: Few-shot graph classification performance (%) of Morpher with DistilBERT (Sanh et al., 2019) as language encoder. Other experiment settings are identical to the main experiment.
@@
B.4 Experiment with GNNs trained using GraphMAE and MVGRL
In the main pages, we used GraphCL and SimGRACE to show that Morpher achieves better performance given a pre-trained GNN. Additionally, to further verify the robustness of Morpher over the pre-train method, we conducted experiments on the pre-trained GNNs using GraphMAE (Hou et al., 2022) and MVGRL (Hassani and Ahmadi, 2020). We use GCN as the GNN backbone and RoBERTa as the LLM encoder, and the results are reported in Table 9 and Table 10.
Table 9: Few-shot graph classification performance (%) of Morpher with the GNN pre-trained by GraphMAE (Hou et al., 2022). Other experiment settings are identical to the main experiment.
@@ B.4 Experiment with GNNs trained using GraphMAE and MVGRL
Table 10: Few-shot graph classification performance (%) of Morpher with the GNN pre-trained by MVGRL (Hassani and Ahmadi, 2020). Other experiment settings are identical to the main experiment.